# Positional Encoding Helps Recurrent Neural Networks Handle a Large Vocabulary

### v3

@@ 1 Introduction
Positional encoding is a high-dimensional representation of time indices on input data (Gehring et al.,, 2017). For instance, its most basic and popular implementation utilizes sinusoidal waves of various predefined frequencies (Vaswani et al.,, 2017); each time step indexed by $t$ is encoded by a vector of $\sin(\omega_{i}t)$ and $\cos(\omega_{i}t)$ with a bundle of frequencies $\omega_{i}$ ($i=1,...,D/2$, where $D$ denotes the total dimensionality of the vector). Positional encoding “timestamps” input tokens by addition/concatenation of these vectors to the corresponding input embeddings. Unlike RNNs, the time representation via positional encoding is invariant to input values (i.e., autonomous) until they are processed together by a network.
@@
Figure 2: Token-wise accuracy of the reverse-ordering task performed by RNNs with and without positional encoding (labeled as “Position-Encoded” and “Vanilla” respectively). The input length was fixed at $64$. The error bars represent the 95% confidence interval estimated from 10,000 bootstrapped samples of five training-test trials with different random seeds. Each of the five trials held out 1024 random sequences (= 65,536 tokens) for computing the test accuracy.
The contributions of this study are summarized as follows:
  • Difficulties in training RNNs on a large vocabulary are demonstrated through systematically designed benchmark tasks. This problem has not been identified in the previous studies—or at most has received little attention from the research community—despite its relevance to empirical applications.
2 Results
2.1 Key Findings
@@
Effects of positional encoding on RNNs were investigated through the reverse-ordering task; [^1] Investigations of additional tasks are reported in Appendix A. given a sequence of random integers, two variants of RNNs, gated recurrent unit (GRU; Cho et al.,, 2014) and long short-term memory (LSTM; Hochreiter and Schmidhuber,, 1997), were trained to reconstruct them in reverse order (e.g. $8,29,2,11\mapsto 11,2,29,8$; Fig. 1). This study adopted the canonical sinusoidal positional encoding (Vaswani et al.,, 2017), [^2] Alternative implementations of positional encoding are also discussed in Appendix C. which was concatenated (rather than added) with the input embeddings (Jaegle et al.,, 2021).
As a result, positional encoding improved the ability of RNNs to handle a larger vocabulary in the reverse-ordering task (Fig. 2). The position-encoded GRU and LSTM successfully reversed the input sequences of 64 integers [^3] See Appendix B for a discussion of robustness to variable input length. drawn uniformly at random from the vocabularies of size 32–256 and 256–16,384, respectively, achieving the token-wise accuracy above 95%. By contrast, the performance of the vanilla models without positional encoding degraded as the vocabulary size increased.
2.2 Frequency Matters
The most apparent consequence of the increased vocabulary size was the reduced chance of observing individual vocabulary items (e.g., $1/256\to 1/16{,}384$). Accordingly, additional experiments were conducted with non-uniformly distributed tokens to investigate the relation between their frequency vs. RNN performance. Specifically, the input vocabulary was evenly divided into Frequent and Rare groups, and the Frequent tokens had three times the probability of the Rare tokens; that is, the probability of each Frequent token was $3/4\times 2/K$ (where $K$ denotes the total vocabulary size and was set to 64 and 1024 for GRU and LSTM, respectively) whilist the probability of each Rare token was $1/4\times 2/K$.
The experiment revealed that it was the disturbant tokens whose frequency significantly impacted the performance of the vanilla RNNs (Fig. 3). On the one hand, the Rare targets were successfully retrieved as long as they were surrounded by the Frequent disturbants. On the other hand, the vanilla GRU struggled to recover the Frequent targets when the other input tokens were filled with the Rare disturbants. The LSTM performance was also degraded, especially when the targets were located in the first quarter of the input sequence ($1\leq t\leq 16$).
Figure 3: Token-wise accuracy of the reverse-ordering task performed by RNNs with and without positional encoding (labeled as “Position-Encoded” and “Vanilla” respectively). The vocabulary was evenly split into Frequent and Rare groups (32+32 for GRU and 512+512 for LSTM), and the former was sampled three times more frequently than the latter. The input length was fixed at $64$. The error bars represent the 95% confidence interval estimated from 10,000 bootstrapped samples of five training-test trials with different random seeds. Each of the five trials held out 4096 test sequences (= 262,144 tokens), consisting of a single “target” token (frequent or rare) surrounded by 63 “disturbants” (all frequent or all rare). That is, sixteen test sequences were held out for each condition (frequent/rare target $\times$ frequent/rare disturbants $\times$ target positions).
By contrast, the position-encoded RNNs exhibited robustness to the frequency of the target and disturbant tokens. They achieved nearly perfect accuracies in most cases, except when the GRU processed the fully Rare data whose target was located in the first half of the sequence ($1\leq t\leq 32$).
2.3 Analysis of Gradient Stability
[…]
Figure 4: Schematic illustration of the analysis of gradient stability. The RNN trained on the reverse-ordering task processed a pair of input sequences that shared the initial token ($t=1$; blue) but differed in the rest ($2\leq t\leq L$; referred to as “Disturbant A/B”, and colored in orange/purple). For each dimension $i$ of the final RNN output $h_{2L,i}$ at time $2L$, the distant gradient $\left(\frac{\partial h_{2L,i}}{\partial z_{1,1}},\dots,\frac{\partial h_{2L,i}}{\partial z_{1,(2)D}}\right)^{T}$ at the first updated latent state $\vec{\mathbf{z}}_{1}$ ($=\vec{\mathbf{h}}_{1}$ in GRU; $=$ concatenation of the hidden and cell states in LSTM, doubling the total dimensionality to $2D$) was computed per input sequence via backpropagation through time (dashed lines). The gradient stability was defined by the dot-product similarity of the paired gradients normalized over the output dimensions by the coefficients $\alpha_{i}^{(s)}(s\in\{A,B\})$, whose definition is provided in Materials and Methods.
To delve deeper into the influence of token frequency on the RNN performance, the gradients of the RNN latent states were scrutinized. In the analysis, pairs of input sequences were processed by the RNNs trained on the dual-frequency vocabulary (comprising Frequent and Rare items; Fig. 4). Each pair of sequences shared the same initial token ($t=1$; “target”) but varied in the subsequent tokens ($2\leq t\leq L$; “disturbants”). Then, gradients were computed for the distant mapping between the first and last updated states (i.e., at time $t=1$ and $2L$) of the RNNs using backpropagation through time. The stability of RNN learning was assessed by measuring the dot-product similarity of the gradients between the paired input sequences (after normalization over output dimensions; see Materials and Methods for the formal definition).
Figure 5: Gradient stability of RNNs trained on the reverse-ordering task with and without positional encoding (labeled as “Position-Encoded” and “Vanilla” respectively). The stability was defined by the dot-product similarity of latent-to-latent gradients after normalization over the output dimensions, conditioned on two input sequences sharing the initial “target” token (whose Frequent vs. Rare distinction is represented by the line color), followed by Frequent or Rare disturbants (represented by the solid vs. dashed lines). The total input length was $1+63=64$. The average similarity over 1024 input pairs times five trials is plotted for every 5000 training iterations.
3 Discussion
3.1 Preliminaries: Theoretical and Empirical Computational Power of (Vanilla) RNNs
In practice, however, RNN weights are bounded by finite precision and must be optimized based on finite observations of data. These settings impose limitations on the empirical capabilities of RNNs (Chen et al.,, 2018; Weiss et al.,, 2018). For example, empirical RNNs cannot store infinitely many observations in their memory, or state vector(s), and the memorized information decays over time. This latter problem with the memory duration has attracted the interest of researchers, leading to extensive exploration of RNN architectures for a longer-lasting memory (Hochreiter and Schmidhuber,, 1997; Arjovsky et al.,, 2016; Neil et al.,, 2016; Chang et al.,, 2017; Jing et al.,, 2017, 2019; Voelker et al.,, 2019; Gu et al.,, 2020).
3.2 Difficulties in Handling a Large Vocabulary
This study examined the RNN gradients and identified their destabilization when processing low-frequency tokens. Specifically, inputs that do not contribute to gradient-based optimization at a target time step (e.g., tokens at $2\leq t\leq L$ upon the retrieval of the initial token at $t=2L$ in the reverse-ordering task) were found to be detrimental.
In most cases of time series processing, data points carring crucial information for specific time steps become irrelevant otherwise. Consequently, each token exhibits a dual nature—both crucial and noisy—throughout the task, and processing rare tokens is particularly challenging presumably because they are irrelevant at most of the time while making a large impact on the learning through the greater loss to compensate for their fewer learning opportunities. Dealing with such “unignorable noise” presents a pervasive challenge for RNNs.
Although low-frequency tokens destabilize the gradient-based learning of RNNs, the present study also discovered that this issue can be alleviated by positional encoding. This enhancement of RNNs by positional encoding is remarkable because RNNs were specifically designed to process time series data on their own; thus, unlike Transformers, they are presumed to function without relying on an “external clock” (Siegelmann and Sontag,, 1992, 1994, 1995; Siegelmann,, 1996, 1999). Consequently, position-encoded RNNs have remained largely unexplored, despite the widespread applications of positional encoding in various areas of machine learning, including the representation of the elapsed time in diffusion processes (Ho et al.,, 2020; Song et al.,, 2021) and even the spatial information of three-dimensional meshes and point clouds (Mildenhall et al.,, 2020; Jun and Nichol,, 2023).
To the best of the author’s knowledge, only two studies have previously explored position-encoded RNNs. [^5] In other studies, positional encoding and RNNs have served as submodules within more complex models, typically in conjunction with an attention mechanism (Kim et al.,, 2018; Song et al.,, 2020). Karanikolos and Refanidis, (2019) reported that a position-encoded LSTM outperformed a vanilla LSTM as well as a shallow Transformer (with four layers) in text summarization. In another study—predating the proposal of sinusoidal positional encoding in the deep learning community—Vincent-Lamarre et al., (2016) showed that oscillatory signals at random frequencies improved the performance of a random RNN (i.e., reservoir computer) in a timing task, assessing the model’s memory duration by its ability to produce a (smoothed) output pulse after a specified time interval from an onset signal. The findings of the present study—the improvement in the manageable vocabulary size due to the enhanced gradient stability—broaden the currently limited understanding of the impact of positional encoding on RNNs.
4 Materials and Methods
@@ 4.1 Model Architecture
The investigations in this study were all based on single-layer GRU and LSTM combined with input-embedding and output-projection layers (Fig. 1). Each of the integers in the input sequences was first embedded and concatenated with the positional encoding, and then fed to the RNN. After reading the entire input sequence, the network received a command to return the output. This command was represented in the form of a time-invariant learnable vector, and fed to the RNN in place of the input embedding (cf. Arjovsky et al.,, 2016). The outputs from the RNN module were linearly projected into the classification logits, whose cross-entropy loss against the target sequence was used to optimize the entire network. Model predictions in the testing phase were defined by the argmax of these logits for each time step.
This study adopted the canonical sinusoidal positional encoding designed for Transformers (Vaswani et al.,, 2017, see Appendix C for discussions on alternative implementations); specifically, each time step $t$ was encoded by the $D_{pos}$-dimensional vector, $(PE_{t,1},\dots,PE_{t,D_{pos}})^{T}$, defined as follows: [^6] The adjustment of the time and frequency indices (“$t-1$” and “$i-1$”) in Eqs. 1 and 2 aligns the 1-based indexing in this paper (adopted for better readability) with the 0-based indexing in Python.
@@ 4.2 Implementation Details
Across the experiments, the dimensionality of the hidden layer of the RNNs was set to 512. The embedding of the input integers and the memory cell of the LSTM also had the same dimensionality of 512.
@@
It is of note that neither extra training iterations nor greater batch sizes improved the performance of the vanilla models.
@@
All the experiments were implemented in PyTorch (ver. 2.1.1; Paszke et al.,, 2017, 2019) and each training-test trial was executed on a single NVIDIA A100 GPU (with 80GB VRAM) hosted by the Academic Center for Computing and Media Studies, The source code will become available in the open repository upon the publishment of this study.
4.3 Metric of Gradient Stability

### v5

@@ 1 Introduction
Positional encoding is a high-dimensional representation of time indices on input data (Gehring et al.,, 2017). For instance, its most basic and popular implementation utilizes sinusoidal waves of various predefined frequencies (Vaswani et al.,, 2017). Positional encoding “timestamps” input tokens by addition/concatenation of these vectors to the corresponding input embeddings. Unlike RNNs, the time representation via positional encoding is invariant to input values (i.e., autonomous) until they are processed together by a network.
@@
The contributions of this study are summarized as follows:
  • Difficulties in training RNNs on a large vocabulary are demonstrated through systematically designed benchmark tasks. This problem has not been identified in the previous studies—or at most has received little attention from the research community—despite its potential relevance to empirical applications.
2 Related Studies
2.1 Theoretical and Empirical Computational Power of (Vanilla) RNNs
In practice, however, RNN weights are bounded by finite precision and must be optimized based on finite observations of data. These settings impose limitations on the empirical capabilities of RNNs (Chen et al.,, 2018; Weiss et al.,, 2018). For example, empirical RNNs cannot store infinitely many observations in their memory, or state vector(s), and the memorized information decays over time. This latter problem with the memory duration has attracted the interest of researchers, leading to extensive exploration of RNN architectures for a longer-lasting memory (Hochreiter and Schmidhuber,, 1997; Arjovsky et al.,, 2016; Neil et al.,, 2016; Chang et al.,, 2017; Jing et al.,, 2017, 2019).
More recently, research into prolonged memory retention has shifted towards continuous-time models (Voelker et al.,, 2019; Gu et al.,, 2020). Instead of representing the memory of an input sequence through discrete-time modifications of a latent state, these models approximate the input history by a linear combination of orthogonal polynomials in continuous-time space. Consequently, the coefficients of these polynomials yield a finite-dimensional representation of the input sequence (termed the High-Order Polynomial Projection Operator, or HiPPO) and the dynamics of these coefficients can be articulated through an ordinary differential equation (ODE; Gu et al.,, 2020). This concept of continuous-time memory representation has subsequently been extended to neural state-space models by replacing the fixed state matrix in the ODE—corresponding to a manually selected basis of polynomials—with a learnable one, while constraining its structure to a diagonal (plus a row-rank) matrix (Gu et al.,, 2021; Gu et al., 2022a, ; Gu et al., 2022b, ; Gu and Dao,, 2024). Most strikingly, with additional refinements, the latest state-space model has achieved superior language modeling performance, rivaling that of Transformer-based models.
2.2 Positional Encoding
Positional encoding is a high-dimensional representation of temporal structures in input data (Gehring et al.,, 2017). The primary demand for this representation scheme comes from Transformers. Unlike RNNs, Transformers lack an inherent mechanism for representing input arrangements. Accordingly, input tokens to a Transformer are “time-stamped” via addition/concatenation of a position-encoding vector.
In the original implementation of Transformer, token positions were encoded by sinusoidal waves of various predefined frequencies (Vaswani et al.,, 2017). While this original encoding scheme is powerful enough for a wide range of tasks, researchers have explored for other possibilities as well. For example, the well-known BERT pretraining for natural language processing employed learnable embeddings to encode token positions (Devlin et al.,, 2019). Some studies also suggested that the combination of the sinusoidal and learnable encoding improves model performance (Dai et al.,, 2019). There is also an option to encode the distance between tokens rather than the elapsed time from the sequence onset (Shaw et al.,, 2018; Dai et al.,, 2019).
Besides Transformers, positional encoding is also used to represent the elapsed time in diffusion processes (Ho et al.,, 2020; Song et al.,, 2021). Moreover, the effectiveness of positional encoding is not limited to temporal information; previous studies in three-dimensional mesh/point-cloud modeling report that sinusoidal transformation of spatial data improves model performance compared to the raw coordinate representation (Mildenhall et al.,, 2020; Jun and Nichol,, 2023).
Despite the widespread use of positional encoding in various areas of machine learning today, its applications to pure RNNs remain largely unexplored. To the best of the author’s knowledge, only two studies have previously explored position-encoded RNNs. [^2] In other studies, positional encoding and RNNs have served as submodules within more complex models, typically in conjunction with an attention mechanism (Kim et al.,, 2018; Song et al.,, 2020). Karanikolos and Refanidis, (2019) reported that a position-encoded LSTM outperformed a vanilla LSTM as well as a shallow Transformer (with four layers) in text summarization. In another study—predating the proposal of sinusoidal positional encoding in the deep learning community—Vincent-Lamarre et al., (2016) showed that oscillatory signals at random frequencies improved the performance of a random RNN (i.e., reservoir computer) in a timing task, assessing the model’s memory duration by its ability to produce a (smoothed) output pulse after a specified time interval from an onset signal.
Similarly, the time index in time series data has rarely been utilized directly by RNNs, presumably because of its “redundancy” along the functionality of RNNs. As an exception, Neil et al., (2016) proposed a periodic gating mechanism for updating the state and memory cell of LSTM. This periodic gating was scheduled according to a triangular wave interspersed with a plateau at the floor value ($=0.0$; the frequency, phase, and the duration of the wave phase were learnable parameters).
3 Methods
@@
3.1 Task
@@
Effects of positional encoding on RNNs were investigated through the reverse-ordering task; [^3] Investigations of additional tasks are reported in Appendix A. given a sequence of random integers, RNNs were trained to reconstruct them in reverse order (e.g. $8,29,2,11\mapsto 11,2,29,8$; Fig. 1).
@@ 3.2 Model Architecture
The investigations in this study were based on single-layer gated recurrent unit (GRU; Cho et al.,, 2014) and long short-term memory (LSTM; Hochreiter and Schmidhuber,, 1997) combined with input-embedding and output-projection layers (Fig. 1). Besides these RNNs, a neural state-space model, S4D (i.e., S4 with a diagonal state matrix; Gu et al., 2022a, ), was also investigated. Each of the integers in the input sequences was first embedded and concatenated with the positional encoding, and then fed to the RNN/S4D. [^4] The concatenation of the positional encoding and input embeddings inflates the number of learnable parameters in the input-to-hidden weights of RNNs in comparison to the vanilla models. However, this additional parameterization is designed not to impact the learning of input embeddings. Appendix C experimentally demonstrates that merely increasing the model size does not enhance the performance of RNNs. After reading the entire input sequence, the network received a command to return the output. This command was represented in the form of a time-invariant learnable vector, and fed to the RNN in place of the input embedding (cf. Arjovsky et al.,, 2016). The outputs from the RNN/S4D module were linearly projected into the classification logits, whose cross-entropy loss against the target sequence was used to optimize the entire network. Model predictions in the testing phase were defined by the argmax of these logits for each time step.
This study adopted the canonical sinusoidal positional encoding designed for Transformers (Vaswani et al.,, 2017, see Appendix D for discussions on alternative implementations); specifically, each time step $t$ was encoded by the $D_{pos}$-dimensional vector, $(PE_{t,1},\dots,PE_{t,D_{pos}})^{T}$, defined as follows: [^5] The adjustment of the time and frequency indices (“$t-1$” and “$i-1$”) in Eqs. 1 and 2 aligns the 1-based indexing in this paper (adopted for better readability) with the 0-based indexing in Python.
@@ 3.3 Implementation Details
Across the experiments, the dimensionality of the hidden layer of the RNNs was set to 512. The embedding of the input integers and the memory cell of the LSTM also had the same dimensionality of 512. Similarly, the hidden dimensionality of S4D was set to 512, while its state size (or the order of the Legendre polynomials) was maintained at the default value of 64.666The S4D was implemented based on the official code provided in https://github.com/state-spaces/s4/blob/main/models/s4/s4d.py.
@@
All the experiments were implemented in PyTorch (ver. 2.1.1; Paszke et al.,, 2017, 2019) and each training-test trial was executed on a single NVIDIA A100 GPU (with 80GB VRAM) hosted by the Academic Center for Computing and Media Studies, Kyoto University. The source code is available in https://github.com/tkc-morita/position-encoded_rnn.
4 Results
4.1 Key Findings
Figure 2: Token-wise accuracy (left) and sequence-wise reconstruction errors (right) of the reverse-ordering task performed by GRU/LSTM/S4D with and without positional encoding (labeled as “Position-Encoded” and “Vanilla” respectively). The input length was fixed at $64$. The error bars represent the 95% confidence interval estimated from 10,000 bootstrapped samples of five training-test trials with different random seeds. Each of the five trials held out 1024 random sequences (= 65,536 tokens) for computing the test accuracy.
As a result, positional encoding improved the ability of RNNs to handle a larger vocabulary in the reverse-ordering task (Fig. 2). The position-encoded GRU and LSTM successfully reversed the input sequences of 64 integers [^7] See Appendix B for a discussion of robustness to variable input length. drawn uniformly at random from the vocabularies of size 32–256 and 256–16,384, respectively, achieving the token-wise accuracy above 95%. By contrast, the performance of the vanilla models without positional encoding degraded as the vocabulary size increased. Similarly, positional encoding enhanced the capacity of S4D to handle large vocabularies. These improvements are also evident in the reduced sequence-wise reconstruction errors, measured by the Damerau-Levenshtein distance. It is of note that neither extra training iterations nor greater batch sizes improved the performance of the vanilla models.
4.2 Frequency Matters
The most apparent consequence of the increased vocabulary size was the reduced chance of observing individual vocabulary items (e.g., $1/256\to 1/16{,}384$). Accordingly, additional experiments were conducted with non-uniformly distributed tokens to investigate the relation between their frequency vs. RNN performance. Specifically, the input vocabulary was evenly divided into Frequent and Rare groups, and the Frequent tokens had three times the probability of the Rare tokens; that is, the probability of each Frequent token was $7/8\times 2/K$ (where $K$ denotes the total vocabulary size and was set to 64, 1024, 2048 for GRU, LSTM, and S4D respectively) whilist the probability of each Rare token was $1/8\times 2/K$.
Figure 3: Token-wise accuracy of the reverse-ordering task performed by GRU/LSTM/S4D with and without positional encoding (labeled as “Position-Encoded” and “Vanilla” respectively). The vocabulary was evenly split into Frequent and Rare groups (32+32 for GRU, 512+512 for LSTM, and 1024+1024 for S4D), and the former was sampled three times more frequently than the latter. The input length was fixed at $64$. The error bars represent the 95% confidence interval estimated from 10,000 bootstrapped samples of five training-test trials with different random seeds. Each of the five trials held out 4096 test sequences (= 262,144 tokens), consisting of a single “target” token (frequent or rare) surrounded by 63 “disturbants” (all frequent or all rare). That is, sixteen test sequences were held out for each condition (frequent/rare target $\times$ frequent/rare disturbants $\times$ target positions).
The experiment revealed that it was the disturbant tokens whose frequency significantly impacted the performance of the vanilla RNNs and S4D (Fig. 3). On the one hand, the Rare targets were successfully retrieved as long as they were surrounded by the Frequent disturbants. On the other hand, the vanilla GRU struggled to recover the Frequent targets when the other input tokens were filled with the Rare disturbants. The LSTM performance was also degraded, especially when the targets were positioned in the first quarter of the input sequence ($1\leq t\leq 16$). Similarly, the Rare disturbants were detrimental to the S4D; unlike the RNNs, however, the accuracy was worst when the targets were located in the middle of the input sequences ($17\leq t\leq 32$).
By contrast, the position-encoded RNNs exhibited robustness to the frequency of the target and disturbant tokens. They achieved nearly perfect accuracies in most cases, except when the GRU processed the fully Rare data whose target was located in the first half of the sequence ($1\leq t\leq 32$). Likewise, positional encoding enhanced the resilience of the S4D against the influence of Rare disturbants.
4.3 Analysis of Gradient Stability
[…]
Figure 4: Schematic illustration of the analysis of gradient stability. The RNN trained on the reverse-ordering task processed a pair of input sequences that shared the initial token ($t=1$; blue) but differed in the rest ($2\leq t\leq L$; referred to as “Disturbant A/B”, and colored in orange/purple). For each dimension $i$ of the final RNN output $h_{2L,i}$ at time $2L$, the distant gradient $\left(\frac{\partial h_{2L,i}}{\partial z_{1,1}},\dots,\frac{\partial h_{2L,i}}{\partial z_{1,(2)D}}\right)^{T}$ at the first updated latent state $\vec{\mathbf{z}}_{1}$ ($=\vec{\mathbf{h}}_{1}$ in GRU; $=$ concatenation of the hidden and cell states in LSTM, doubling the total dimensionality to $2D$) was computed per input sequence via backpropagation through time (dashed lines). The gradient stability was defined by the dot-product similarity of the paired gradients normalized over the output dimensions by the coefficients $\alpha_{i}^{(s)}(s\in\{A,B\})$, whose definition is provided in Eq. 4.
To delve deeper into the influence of token frequency on the RNN performance, the gradients of the RNN latent states were scrutinized. In the analysis, pairs of input sequences were processed by the RNNs trained on the dual-frequency vocabulary (comprising Frequent and Rare items; Fig. 4). Each pair of sequences shared the same initial token ($t=1$; “target”) but varied in the subsequent tokens ($2\leq t\leq L$; “disturbants”). Then, gradients were computed for the distant mapping between the first and last updated states (i.e., at time $t=1$ and $2L$) of the RNNs using backpropagation through time. The stability of RNN learning was assessed by measuring the dot-product similarity of the gradients between the paired input sequences (after normalization over output dimensions).